In 4 years a mother will be five times as old as her daughter at present the mother is 9 times as old as the daughter how old are the mother and the daughter today
Answer by lwsshak3(11628)

In 4 years a mother will be five times as old as her daughter at present the mother is 9 times as old as the daughter how old are the mother and the daughter today
let x=daughter's present age
9x- mother's present age
in 4 yrs:
x+4=daughter's age
9x+4= mother's age
9x+4=5(x+4
9x+4=5x+20
4x=16
x=4
9x=36
daughter's present age=4
mother's present age=36